# Doublons Bloch oscillations in the mass-imbalanced extended Fermi-Hubbard model

### v1

@@ III.2 Free doublons dynamics in an interacting system
Now we investigate the BOs at the resonance point of system parameters, which is $V_{\mathrm{eff}}=V-(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U=0$. We are interested in the multi-doublon dynamics. Obviously, the released doublons initial state acts like free hard-core bosons according to the doublons effective Hamiltonian in Eq. (5), but with renormalized NN hopping amplitude $J_{\mathrm{eff}}$. In Figs. 4(a) and (c), we present the numerical results of time evolution for the initial state in Eq. (15) with different NN interaction but without the tilted potential. It is shown that the deviation from the resonance condition $V=(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$ obstructs the transport of doublons. This can be demonstrated by the following correlation function that characterize the mean-square displacement (MSD) [78] between the doublons
@@
The numerical results of the MSDs for these two cases are presented in Fig. 5. For the resonant case, we can observe the characteristic of ballistic transport, i.e., $\overline{\mathcal{D}^{2}}(t)\sim t^{2}$, which means the doublons spread without scattering.
@@
where the doubling of the tilted potential comes form the double occupancy of fermions per site. This yields the WS ladder for free doublons. The period of the BOs would be a half of that of free particle in lattice. Numerical result in Fig. 4(d) shows the free-doublon BOs at the resonance point. From Figs. 4(a), (b) and the corresponding MSDs in Fig. 5, we can see that a minor deviation from the parameter $V=(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$ induces decoherence of the diffusion and BO, highlighting the effect of NN interaction. In Figs. 4(e) and (f), we present the results for the initial state with $4$ doublons ($8$ fermions)
@@
corresponding to Figs. 4(c) and (d), but with a smaller system size $L=20$. The larger tilted potential in Fig. 4(f) is taken to constrain the amplitude of BO, avoiding the boundary effect. The diffusion behavior and BO are similar with the cases of $2$ doublons. The oscillating periods for the breathing modes are $T=2\pi/(2F)\approx 628.32$ and $314.16$ for the cases in Figs. 4(d) and (f), respectively.
@@ IV Numerical results of two-dimensional system
Taking two NN doublons localized in the center of the lattice as an initial state, Figs. 6(a) and (b) show the numerical results of doublon density at different time for zero and nonzero tilted potential $\mathbf{F}$, respectively. As expected, Fig. 6(a) shows that the doublons manifest the spreading dynamics resembling the free particles in a square lattice. While the doublons in Fig. 6(b) undergo the BOs, with periods $T_{x}=2\pi/(2F_{x})=104.72$ and $T_{y}=2\pi/(2F_{y})=52.36$ in $x$ and $y$ directions, respectively.
@@ V Conclusion
We have studied multi-doublon BOs in the extended Fermi-Hubbard model with imbalance mass. The doublons effective Hamiltonian is given to elucidate the mechanisms of two types of doublons BOs, i.e., the bound-doublon and free-doublon BOs. The interaction between doublons stems not only from the NN interaction between fermions but also from the hopping and on-site interaction. Therefore, there is a resonance point in the system parameters that results in the free-doublon dynamics. It is shown that the spin-up and spin-down fermions are always well bounded as doublons and exhibit two types of BOs dynamics, despite of the imbalance of their mass. Numerical results of the time evolution of doublon density and MSD demonstrate our conclusions. The numerical simulations of the system on a square lattice suggest that the analysis can be applied to higher dimensions. Our findings highlight the role of NN interaction in the nonequilibrium dynamic properties, as well as pave the way for the future explorations of BOs in strongly correlated systems.
APPENDIX Derivation of the doublons effective Hamiltonian
@@
$+\hat{n}_{j,\bar{\sigma}}^{\dagger}\hat{c}_{j,\sigma}^{\dagger}\hat{c}_{i,\sigma}(1-\hat{n}_{i,\bar{\sigma}})],$
@@
where $\hat{S}$ requires to be a Hermitian operator, so that the above transformation is unitary. We firstly wish to eliminate the term $\hat{H}_{J}^{+}+\hat{H}_{J}^{-},$ which can be done by selecting the $\hat{S}$ operator such that $\hat{H}_{J}^{+}+\hat{H}_{J}^{-}+\mathrm{i}\left[\hat{S},\hat{H}_{U}\right]=0.$ Note that $\left[\hat{H}_{J}^{\pm},\hat{H}_{U}\right]=\mp U\hat{H}_{J}^{\pm}$, which suggests the form
$\hat{S}=-\frac{\mathrm{i}}{U}\left(\hat{H}_{J}^{+}-\hat{H}_{J}^{-}\right).$ (A5)
@@
$\hat{H}_{\mathrm{eff}}$ $=$ $\hat{H}_{U}+\hat{H}_{V}+\hat{H}_{J}^{0}+\mathrm{i}\left[\hat{S},\hat{H}_{V}\right]+\frac{1}{U}\left[\hat{H}_{J}^{+},\hat{H}_{J}^{-}\right]$ (A6)
$+\mathrm{i}\left[\hat{S},\hat{H}_{J}^{0}\right]+O\left(J^{2}V/U^{2}\right)+O\left(J^{3}/U^{2}\right).$
Here we set $O(J_{\uparrow})=O(J_{\downarrow})=O(J).$ To eliminate the term $\left[\hat{S},\hat{H}_{J}^{0}\right],$ we note that $\hat{S}=O\left(J/U\right),$ thus we can consider a second order correction of $\hat{S}$ , that is
$\hat{S}\rightarrow\hat{S}+\hat{S}^{(2)},$ (A7)
with $\hat{S}^{(2)}=O\left(J^{2}/U^{2}\right)$ being determined by the operator equation $\mathrm{i}\left[\hat{S}^{(2)},\hat{H}_{U}\right]+\left[\hat{H}_{J}^{+}-\hat{H}_{J}^{-},\hat{H}_{J}^{0}\right]/U=0.$
Due to a large $\left|U\right|$, the doublons states are well separated from others with the energy gaps of order $\left|U\right|$. Then the Hamiltonian can be safely projected into the doublons subspace up to order $J^{2}/U$. The effective Hamiltonian in Eq. (A6) can be simplified by taking the following consideration into account: (i) Term $[\hat{S},\hat{H}_{V}]$ can be omitted, since it connects the doublons subspace to others. (ii) We neglect term $[\hat{S},[\hat{S},\hat{H}_{V}]]=O\left(J^{2}V/U^{2}\right)$, since it is higher than order $J^{2}/U$ and involves multi-site processes. (iii) Term $\hat{H}_{J}^{0}$ describing the hopping processes between singly occupied sites and empty (or doubly occupied) sites, also vanishes in the doublons subspace.
@@
where $\sum_{\left\langle i,j,m\right\rangle}$ denotes the summation over trimers with site $j$ being a nearest neighbor of both sites $i$ and $m$. It can be checked that this three-site processes vanish in the doublons subspace, and the two-site processes remain. Direct calculation shows that

### v3

@@ III.2 Free doublons dynamics in an interacting system
Now we investigate the BOs at the resonance point of system parameters, which is $V_{\mathrm{eff}}=V-(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U=0$. We are interested in the multi-doublon dynamics. Obviously, the released doublons initial state acts like free hard-core bosons according to the doublons effective Hamiltonian in Eq. (5), but with renormalized NN hopping amplitude $J_{\mathrm{eff}}$. Apart from the resonance nature, it is also worth noting that the doublons BOs remain, despite the presence of hard-core constraint. In Figs. 4(a) and (c), we present the numerical results of time evolution for the initial state in Eq. (15) with different NN interaction but without the tilted potential. It is shown that the deviation from the resonance condition $V=(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$ obstructs the transport of doublons. This can be demonstrated by the following correlation function that characterize the MSD [87] between the doublons
@@
The definition of MSD involves the doublon density correlator, which can be measured through the quantum gas microscopy [84]. The numerical results of the MSDs for these two cases are presented in Fig. 5. For the resonant case, we can observe the characteristic of ballistic transport, i.e., $\overline{\mathcal{D}^{2}}(t)\sim t^{2}$, which means the doublons spread without scattering.
@@
where the doubling of the tilted potential comes form the double occupancy of fermions per site. This yields the WS ladder for free doublons. The period of the BOs would be a half of that of free particle in lattice. The numerical result in Fig. 4(d) shows the free-doublon BOs at the resonance point. From Figs. 4(a), 4(b) and the corresponding MSDs in Fig. 5, we can see that a minor deviation from the parameter $V=(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$ induces decoherence of the diffusion and BO, highlighting the effect of NN interaction. This can be grasped from the values of the effective parameters, which dominate the doublons dynamics. Although the deviation of $V$ is minor in comparison to $J_{\sigma}$ and $U$, the deviation of the effective NN interaction $\Delta V_{\textrm{eff}}=0.029$ is comparable with the effective hopping amplitudes $J_{\textrm{eff}}=-0.02$. In Figs. 4(e) and (f), we present the results for the initial state with $4$ doublons ($8$ fermions)
@@
corresponding to Figs. 4(c) and (d), but with a smaller system size $L=20$. The larger tilted potential in Fig. 4(f) is taken to constrain the amplitude of BO, avoiding the boundary effect. The diffusion behavior and BO are similar with the cases of $2$ doublons. The oscillating periods for the breathing modes are $T=2\pi/(2F)\approx 628.32$ and $314.16$ for the cases in Figs. 4(d) and 4(f), respectively.
@@ IV Numerical results of two-dimensional system
Taking two NN doublons localized in the center of the lattice as an initial state, Figs. 6(a) and 6(b) show the numerical results of doublon density at different time for zero and nonzero tilted potential $\mathbf{F}$, respectively. As expected, Fig. 6(a) shows that the doublons manifest the spreading dynamics resembling the free particles in a square lattice. While the doublons in Fig. 6(b) undergo the BOs, with periods $T_{x}=2\pi/(2F_{x})=104.72$ and $T_{y}=2\pi/(2F_{y})=52.36$ in $x$ and $y$ directions, respectively.
@@ V Conclusion
We have studied multi-doublon BOs in the extended Fermi-Hubbard model with imbalance mass. The doublons effective Hamiltonian is given to elucidate the mechanisms of two types of doublons BOs, i.e., the bound-doublon and free-doublon BOs. The interaction between doublons stems not only from the NN interaction between fermions but also from the hopping and on-site interaction. Therefore, there is a resonance point in the system parameters that results in the free-doublon dynamics. It is shown that the spin-up and spin-down fermions are always well bounded as doublons and exhibit two types of BOs dynamics, despite the imbalance of their mass. Numerical results of the time evolution of doublon density and MSD demonstrate our conclusions. The numerical simulations of the system on a square lattice suggest that the analysis can be applied to higher dimensions. Although the presence of mass imbalance is not essential for the observation of doublons BOs, in contrast to the nearest-neighbor interaction, it may facilitates the realization in various experimental platforms. Our findings highlight the role of NN interaction in the nonequilibrium dynamic properties, as well as pave the way for the future explorations of BOs in strongly correlated systems.
Acknowledgements.
This work was supported by Research Fund of Guangdong-HongKong-Macao Joint Laboratory for Intelligent Micro-Nano Optoelectronic Technology (No. 2020B1212030010).
APPENDIX A Derivation of the doublons effective Hamiltonian
@@ Acknowledgements.
$+\hat{n}_{j,\bar{\sigma}}\hat{c}_{j,\sigma}^{\dagger}\hat{c}_{i,\sigma}(1-\hat{n}_{i,\bar{\sigma}})],$
@@
where $\hat{S}$ requires to be a Hermitian operator, so that the above transformation is unitary. We would like to keep the term in $\hat{H}_{\mathrm{eff}}$ up to $O(J^{2}/U)$. This can be done by taking $\hat{S}$ up to second order, that is
$\hat{S}=\hat{S}^{(1)}+\hat{S}^{(2)},$ (A5)
where $\hat{S}^{(1)}=O(J/U)$ and $\hat{S}^{(2)}=O(J^{2}/U^{2})$. Here we set $O(J_{\uparrow})=O(J_{\downarrow})=O(J).$ One can check that $\hat{H}_{U}$ and $\hat{H}_{\lambda}$ satisfy
$\left[\hat{H}_{U},\hat{H}_{J}^{\lambda}\right]=\lambda U\hat{H}_{J}^{\lambda},$ (A6)
where $\lambda=0,\pm$. This commutation relation is useful for the upcoming discussion.
Obviously, the largest term in $\left[\hat{S},\hat{H}\right]$ is $\left[\hat{S},\hat{H}_{U}\right]$, which can be exploited to cancel other terms. We firstly wish to eliminate the term $\hat{H}_{J}^{+}+\hat{H}_{J}^{-},$ which can be done by selecting $\hat{S}^{(1)}$ such that
$\hat{H}_{J}^{+}+\hat{H}_{J}^{-}+\mathrm{i}\left[\hat{S}^{(1)},\hat{H}_{U}\right]=0.$ (A7)
Using Eq. (A6), we can check that
$\hat{S}^{(1)}=-\frac{\mathrm{i}}{U}\left(\hat{H}_{J}^{+}-\hat{H}_{J}^{-}\right)$ (A8)
does satisfy Eq. (A7).
To eliminate the term $\left[\hat{S}^{(1)},\hat{H}_{J}^{0}\right]$, we consider the second order term of $\hat{S}$, which is determined by the operator equation
$\mathrm{i}\left[\hat{S}^{(1)},\hat{H}_{J}^{0}\right]+\mathrm{i}\left[\hat{S}^{(2)},\hat{H}_{U}\right]=0.$ (A9)
Again, using Eq. (A6) and the Jacobi identity, we can check that
$\hat{S}^{(2)}=-\frac{\mathrm{i}}{U^{2}}\left[\hat{H}_{J}^{+}+\hat{H}_{J}^{-},\hat{H}_{J}^{0}\right]$ (A10)
satisfy Eq. (A9). Thus the terms $\left[\hat{S}^{(2)},\hat{H}_{J}^{0}\right]$ and $\left[\hat{S}^{(2)},\hat{H}_{J}^{+}+\hat{H}_{J}^{-}\right]$ are of the order of $J^{3}/U^{2}$, and term $\left[\hat{S}^{(2)},\hat{H}_{V}\right]$ is of the order of $J^{2}V/U^{2}$.
In addition, direct calculations yield
$\textrm{i}\left[\hat{S}^{(1)},\hat{H}_{J}^{+}+\hat{H}_{J}^{-}\right]=\frac{2}{U}\left[\hat{H}_{J}^{+},\hat{H}_{J}^{-}\right],$ (A11)
and
$\frac{\mathrm{i}^{2}}{2}\left[\hat{S}^{(1)},\left[\hat{S}^{(1)},\hat{H}_{U}\right]\right]=-\frac{1}{U}\left[\hat{H}_{J}^{+},\hat{H}_{J}^{-}\right].$ (A12)
Other terms in $\mathrm{i}^{2}\left[\hat{S},\left[\hat{S},\hat{H}\right]\right]/2$ have the order that higher than $J^{2}/U$.
@@
$\hat{H}_{\mathrm{eff}}$ $=$ $\hat{H}_{U}+\hat{H}_{V}+\hat{H}_{J}^{0}+\mathrm{i}\left[\hat{S}^{(1)},\hat{H}_{V}\right]+\frac{1}{U}\left[\hat{H}_{J}^{+},\hat{H}_{J}^{-}\right]$ (A13)
$+O\left(J^{2}V/U^{2}\right)+O\left(J^{3}/U^{2}\right).$
In fact, the energy scale $J^{2}/U$ in Eq. (A13) can be grasped from the second order process of degenerate perturbation theory, resembling Eq. (13). For example, when taking the doublons states as the degenerate basis, and the hopping term as perturbation, the hopping of a doublon is achieved by the hopping of two fermions through the second order process, which includes other states possessing energy gaps from the doublons states of order $U$, as an intermediate process. Thus, this process possesses energy scale $J^{2}/U$. Due to a large $\left|U\right|$, the doublons states are well separated from others with the energy gaps of order $\left|U\right|$. Then the Hamiltonian can be safely projected into the doublons subspace up to order $J^{2}/U$. The effective Hamiltonian in Eq. (A13) can be simplified by taking the following consideration into account: (i) Term $[\hat{S}^{(1)},\hat{H}_{V}]$ can be omitted, since it connects the doublons subspace to others. (ii) Term $\hat{H}_{J}^{0}$ describing the hopping processes between singly occupied sites and empty (or doubly occupied) sites, also vanishes in the doublons subspace.
@@
where $\sum_{\left\langle i,j,m\right\rangle}$ denotes the summation over trimers with site $j$ being a nearest neighbor of both sites $i$ and $m$. It can be checked that these three-site processes vanish in the doublons subspace, and the two-site processes remain. Direct calculation shows that
@@
APPENDIX B Chebyshev expansion method for time evolution
For clarity, here we formulate the implementation of the Chebyshev expansion method for numerical simulations of time evolution $\left|\Psi\left(t\right)\right\rangle=\exp(-\mathrm{i}\hat{\mathcal{H}}t)\left|\Psi\left(0\right)\right\rangle$. The more detailed descriptions for this method can be found in Refs. [81, 82, 83].
To expand the time evolution operator into a finite series of Chebyshev polynomials, we first need to rescale the Hamiltonian
~=^$\widetilde{\mathcal{H}}=\frac{\hat{\mathcal{H}}-b}{a},$ (A17)
fitting the spectrum into the interval $\left[-1,1\right]$. The scaling factors are $a=(E_{\text{max}}{-}E_{\text{min}})/(2-\epsilon)$ and $b=(E_{\text{max}}{+}E_{\text{min}})/2$, where we take a small cutoff $\epsilon=0.01$ to ensure the rescaled spectrum lies inside $\left[-1,1\right]$, and the extremal eigenvalues of the Hamiltonian, $E_{\text{max}}$ and $E_{\text{min}}$, can be efficiently computed by the the implicitly restarted Arnoldi method [90]. Then the time evolution operator can be expanded as a series of Chebyshev polynomials
$\exp(-\mathrm{i}\hat{\mathcal{H}}t)\left|\Psi\left(0\right)\right\rangle$ (A18)
$\approx\exp(-\mathrm{i}bt)\left[J_{0}(at)+2\sum_{n=1}^{N}(-\mathrm{i})^{n}J_{n}(at)T_{n}(\widetilde{\mathcal{H}})\right]\left|\Psi\left(0\right)\right\rangle,$
where $J_{n}(x)$ denotes the first-kind Bessel function of order $n$, and $T_{n}(x)=\cos(n\arccos(x))$ denotes the first-kind Chebyshev polynomials of order $n$. The term $|\nu_{n}\rangle=T_{n}(\widetilde{\mathcal{H}})\left|\Psi\left(0\right)\right\rangle$ in Eq. (A18) can be computed iteratively by the following recurrence relation of the Chebyshev polynomials
$|\nu_{n+1}\rangle=2\widetilde{\mathcal{H}}|\nu_{n}\rangle-|\nu_{n-1}\rangle,$
$|\nu_{1}\rangle=\widetilde{\mathcal{H}}|\nu_{0}\rangle,$ (A19)
where $|\nu_{0}\rangle=\left|\Psi\left(0\right)\right\rangle$. The algorithm requires memory only for a few vectors and for the sparse matrix Hamiltonian. For a large $n$, the Bessel function has the asymptotic behavior
$J_{n}(at)\sim\frac{1}{\sqrt{2\pi n}}\left(\frac{\text{e}at}{2n}\right)^{n},$ (A20)
which decay superexponentially for $n\gg at$, and the series in Eq. (A18) can be truncated with negligible error, i.e., we can choose $N$ such that for $n>N$, the expansion coefficients $J_{n}(at)$ is smaller than a desired accuracy cutoff.
The Chebyshev expansion method relies only on matrix vector multiplications, and thus can exploit the sparsity of the matrix representation of the Hamiltonian to save the computer memory and computation time significantly. In addiction, unlike the differencing method, this method does not accumulate errors, and taking a large time step is possible by increasing $N$. It is especially suitable for the computations of time evolution of quantum state under time-independent Hamiltonian.